OpenStudy (thebookkeeper):

PLEASE HELP ME! AB<--> and BC<--> form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC<-->? x + 3y = 16 2x + y = 12 -7x -5y = -48 7x -5y = 48

OpenStudy (anonymous):

Basically youre trying to find the line perpendicular to AB.

OpenStudy (anonymous):

So your first step is to find the equation of the line from the given coordinates.

OpenStudy (thebookkeeper):

ok how do I do that?

OpenStudy (anonymous):

First find the slope of the line, can you do that on your own?

OpenStudy (thebookkeeper):

isn't it (7,-6)?

OpenStudy (anonymous):

No, slope is expressed as \[\frac{ rise }{ run }\] and can be found using \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (anonymous):

The y_1, y_2, x_1 and x_2 are place holders for the two points you were given. Like: (x_1,y_1) (x_2, y_2)

OpenStudy (anonymous):

Hint: The two you were given was (-3, -1) and (4, 4)

OpenStudy (anonymous):

(-3, -1) (4, 4) (x_1, y_1) (x_2, y_2)

OpenStudy (thebookkeeper):

so the slope is 5/7?

OpenStudy (anonymous):

Yes! Good job. We are not done yet however.

OpenStudy (anonymous):

So the standard form of a line is y = mx + b where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

For the sake of this problem we will ignore b.

OpenStudy (anonymous):

The negitave slope will be perpendicular to our line.

OpenStudy (anonymous):

So -5/7x + b = y

OpenStudy (anonymous):

We move things around: 7x - 5y = -7b

OpenStudy (anonymous):

And only one option's right side looks like this: 7x -5y = 48

OpenStudy (thebookkeeper):

that's not the answer though. I tried it and it said it was wrong